OpenStudy (anonymous):

follow principles of estimation to find an approximate value. round each number so that there is one nonzero digit do not find the exact value 4423 * 4997 thanks in advance

OpenStudy (ybarrap):

4423*4997 is about (4000 + 423)(5000) =(4000*5000 + 423*5000) which is about 20,000,000 + 2,000,000=22,000,000 Compare this to the actual answer of 22,101,731. Not too bad!

OpenStudy (ybarrap):

Note that I used 400 instead of 423 above

OpenStudy (anonymous):

by rounding down to the nearest digit?

OpenStudy (ybarrap):

Yes

OpenStudy (anonymous):

much appreciated thank you nicely explained

OpenStudy (ybarrap):

NP